Find the value of b for which the equation
(ye^2xy + x) dx + (bxe^2xy) dy =0
is exact, and hence solve it for that value of b

b) Find the solution of the Riccati equation
dy/dx =2cos^2x-sin^2x+y^2 / 2cosx
y1(x) = sin x
1
Expert's answer
2020-04-02T13:46:55-0400

a) "(ye^{2xy} + x) dx + (bxe^{2xy}) dy=dU", then "ye^{2xy}+x=\frac{\partial U}{\partial x}" and "bxe^{2xy}=\frac{\partial U}{\partial y}".

We have "\frac{\partial^2 U}{\partial x\partial y}=\frac{\partial^2 U}{\partial y\partial x}", that is "\frac{\partial}{\partial x}(bxe^{2xy})=\frac{\partial}{\partial y}(ye^{2xy}+x)"

"\frac{\partial}{\partial x}(bxe^{2xy})=be^{2xy}+2bxye^{2xy}", "\frac{\partial}{\partial y}(ye^{2xy}+x)=e^{2xy}+2xye^{2xy}" , so "be^{2xy}+2bxye^{2xy}=e^{2xy}+2xye^{2xy}", and we obtain "b=1".

Hence "\frac{\partial U}{\partial y}=xe^{2xy}", so "U=\frac{e^{2xy}}{2}+f(x)".

Then "\frac{\partial U}{\partial x}=ye^{2xy}+f'(x)". Since "\frac{\partial U}{\partial x}=ye^{2xy}+x", we have "f'(x)=x".

Let "f(x)=\frac{x^2}{2}", then "U=\frac{e^{2xy}+x^2}{2}"

So the given equation we can rewrite as "d\left(\frac{e^{2xy}+x^2}{2}\right)=0", general solution of which is "\frac{e^{2xy}+x^2}{2}=\frac{C}{2}" or "e^{2xy}+x^2=C"

b)General solution we will look in form "y=y_1+u=u+\sin x"

We have "y'=u'+\cos x", then

"u'+\cos x=\frac{2\cos^2x-\sin^2x+(u+\sin x)^2}{2cosx}"

After expanding brackets and moving "\cos x" to the right side, we obtain

"u'=\frac{u^2+2u\sin x}{2cosx}" or "\frac{du}{dx}=\frac{u^2+2u\sin x}{2cosx}"

So "2\cos xdu=u^2dx+2u\sin xdx" or "2\cos xdu=u^2dx-2ud(\cos x)"

Since "d(UV)=VdU+UdV" for every differentiable "U" and "V", we have "2d(u\cos x)=u^2dx"

Let "u\cos x=v", then "2dv=\frac{v^2}{\cos^2 x}dx".


So we have "v=0" or "\frac{2dv}{v^2}=\frac{dx}{\cos^2 x}" or "-2d\left(\frac{1}{v}\right)=d(\tan x)" . We obtain the general solution "v=0" or "\tan x+\frac{2}{v}=C".

Reurning to "u", we obtain "u=0" or "\tan x+\frac{2}{u\cos x}=C". And, returning to "y", we obtain "y=\sin x" or "\tan x+\frac{2}{(y-\sin x)\cos x}=C". Expessing "y", we obtain "y=\sin x" or "y=\sin x+\frac{2}{C\cos x-\sin x}"


Answer: a)"b=1", "e^{2xy}+x^2=C", b)"y=\sin x" or "y=\sin x+\frac{2}{C\cos x-\sin x}"